# CounterQuill: Investigating the Potential of Human-AI Collaboration in Online Counterspeech Writing

### v1

@@ 5.2. Study Protocol
Each study session lasted approximately 70 minutes and employed a within-subjects design. Before the study, participants received a brief introductory email about the system and a tutorial video that provided a comprehensive guide on using CounterQuill’s features. During the study, participants experienced two conditions: a 10-minute period of writing counterspeech with ChatGPT, and a 50-minute period of co-writing counterspeech with CounterQuill. The detailed study procedure is described in §5.3.
Figure 9. A diagram outlining the study protocol used. The intervention ordering was counterbalanced between participants and participants engaged with different conditions. In addition, the times labelled in this figure are the average time spent by the participants.
@@ 5.2.1. Conditions
In our user study, we implemented two distinct conditions to evaluate the effectiveness of AI-assisted counterspeech writing, drawing on previous research that used GPT-based systems (Zhang et al., 2023; Reza et al., 2024; Lin et al., 2024) such as GPT-3.5 as baseline:
@@
  • CounterQuill: Participants used the full version of CounterQuill, which includes a learning session (help users understand hatespeech and counterspeech), a brainstorming session (step-by-step counterspeech ideation support), and a writing session (co-writing with CounterQuill for counterspeech).
To minimize potential biases, we used a Latin Square experimental design (Ryan and Morgan, 2007) to ensure a balanced sequence of task topics and conditions for each study session: with half the participants starting with our system and the other half beginning with ChatGPT. To assess the potential impact of the system usage order on participants’ perceptions and the overall study outcomes, we conducted Welch’s Two-Sample t-tests that accounts for unequal variances between groups (Sakai, 2016), as detailed in Tables A2 and A3. The results confirmed that the sequence—whether participants began with ChatGPT or CounterQuill—did not significantly influence their performance or feedback on the system.
Although the primary comparison focuses on the writing session, we recognize that the preceding learning and brainstorming sessions could influence the results, and we have addressed this in the limitations and future work section.
@@ 5.3. Study Procedure
At the beginning of each study session, the experimenters collected informed consent from the participants. For the user study, each participant was presented with 20 hate speech examples and asked to select one of them. The hate speech data was sourced from Ping’s (2024) research (Ping et al., 2024), which included a total of 400 hate speech instances evenly distributed across five hate speech themes: race, gender, sexual orientation, disability, and religion. To ensure a balanced representation of themes, each participant was randomly assigned 20 hate speech examples, with an average of 4 examples per theme. Table A1 (Appendix F) lists the hate speech instances used by each of the 20 participants.
The study coordinators introduced participants to the ChatGPT counterspeech writing session and provided a 3-minute tutorial on how to use CounterQuill, followed by an introduction to a CounterQuill co-writing session. Each participant completed the two writing sessions, with the order counterbalanced between participants to minimize potential biases (Greenwald, 1976; Charness et al., 2012).
In the ChatGPT (Baseline) condition, participants used ChatGPT to write counterspeech for the hateful content. Then they completed a post-study questionnaire including NASA-TLX questions and customized questions (see Appendix B for post-study questionnaire). In the CounterQuill condition, participants went through a learning session to understand hatespeech and counterspeech, a brainstorming session for step-by-step ideation support, and used the CounterQuill co-writing assistant. They then completed a post-study questionnaire. The study ended with a 20-minute semi-structured interview, during which the experimenter asked questions to understand participants’ experiences, feedback on CounterQuill features, and reflections on the two conditions. The purpose of the semi-structured interview was to explore participants’ overall experiences, elicit their specific feedback on CounterQuill and ChatGPT for counterspeech writing, and gather suggestions for improvement.
@@ 6. Result
6.1. Quantitative Analysis
6.1.1. Statistical Analysis of ChatGPT and CounterQuill with NASA TLX
Task Average Time (min)
CounterQuill Learning 8.72 $\pm$ 1.02
CounterQuill Brainstorming 10.32 $\pm$ 1.55
CounterQuill Writing 10.48 $\pm$ 1.71
ChatGPT Writing 9.11 $\pm$ 1.06
Table 3 summarizes the ratings for each dimension of the NASA Task Load Index (TLX) across two conditions: using ChatGPT (Condition A) and using CounterQuill (Condition B). We evaluated the differences between these conditions on each dimension using paired-sample $t$-tests (Mowery, 2011). To further contextualize these findings, Table 2 provides a comparison of the average time spent on counterspeech composition in each condition.
@@ 6.1.1. Statistical Analysis of ChatGPT and CounterQuill with NASA TLX
The analysis of the NASA-TLX dimensions revealed several notable differences between Condition A (ChatGPT) and Condition B (CounterQuill). While Mental Demand showed a slight, non-significant increase from Condition A to B, Physical Demand, Temporal Demand, and Effort all displayed statistically significant increases in Condition B. Performance also significantly improved in Condition B, with a large effect size. Conversely, Frustration was the only dimension that showed a significant decrease in Condition B compared to A.
Figure 10. The violin plot of NASA-TLX score distribution: ChatGPT vs. CounterQull.
The violin plot (Figure 10) visually represents the score distributions across the two conditions, further emphasizing the differences highlighted by the statistical analysis. The plot shows the shifts in the distribution of scores for each dimension, with higher scores for Physical Demand, Temporal Demand, Performance, and Effort in using CounterQuill, and lower scores for Frustration.
In summary, using CounterQuill requires more effort, time, and physical energy compared to using ChatGPT. However, it also leads to significantly better perceived performance outcomes, with participants reporting a 63.66% higher rate with CounterQuill (from NASA TLX ‘Performance’). This higher performance perception is not only due to their overall satisfaction with the counterspeech they co-writed but also their strong performance in the quiz after the learning session. Out of 20 participants, 15 answered all quiz questions correctly, while the remaining five answered 2, 2, 3, 3, and 3 questions correctly, resulting in an average accuracy rate of 91.25%. Additionally, CounterQuill also contributed to a 51.96% reduction in participants’ feelings of frustration compared to ChatGPT (from ‘Frustration’).
6.1.2. Quantitatively Analyze User Perceptions and Experiences with ChatGPT and CounterQuill
Our study further investigated participants’ experiences and perceptions when using ChatGPT and CounterQuill for counterspeech writing. To gain deeper insights, we designed six customized questions (see Appendix B for detailed post-study questionnaire) that focused on specific aspects of the participants’ experience, such as confidence in identifying hate speech, effectiveness of the tools in providing guidance, and satisfaction with the final counter-speech produced. The analysis of the six customized questions revealed several notable findings (Table 4):
Table 4. Comparison of participant perceptions between ChatGPT and CounterQuill for counterspeech writing assistance based on six customized questions. The results of paried $t$-test show statistically significant differences favoring CounterQuill in all categories.
@@ 6.1.2. Quantitatively Analyze User Perceptions and Experiences with ChatGPT and CounterQuill
HS Identification Confidence 4.18$\pm$1.88 6.24$\pm$0.75 -4.280 ¡0.001*** [-3.079, -1.039]
Brainstorming Effectiveness 3.24$\pm$1.44 6.35$\pm$0.70 -8.849 ¡0.001*** [-3.865, -2.371]
Self-Efficacy in CS Writing 3.29$\pm$1.83 6.18$\pm$0.81 -7.356 ¡0.001*** [-3.713, -2.052]
Engagement with AI 3.59$\pm$1.73 6.00$\pm$0.87 -5.734 ¡0.001*** [-3.303, -1.520]
Satisfaction with CS 4.00$\pm$1.58 6.47$\pm$0.51 -7.668 ¡0.001*** [-3.154, -1.788]
Willingness to Post Online 3.53$\pm$1.46 6.29$\pm$0.77 -10.444 ¡0.001*** [-3.326, -2.204]
Figure 11. The violin plot of evaluating ChatGPT and CounterQuill impact on counterspeech writing based on our customized questions.
Our result found statistically significant differences (p ¡ 0.001) between CounterQuill and ChatGPT across various measures. Participants reported higher confidence in identifying online hate speech, found CounterQuill more effective for brainstorming and increasing self-efficacy in writing counterspeech, felt more engaged and motivated when using CounterQuill, were more satisfied with the counterspeech created using CounterQuill, and were more likely to post the co-written counterspeech online when using CounterQuill compared to ChatGPT.
Overall, the violin plot (Figure 11) visually represents the score distributions across the two conditions based on our six customized questions. Participants found CounterQuill more effective in enhancing their ability to identify hate speech, providing brainstorming assistance, increasing self-efficacy, fostering engagement, and leading to satisfaction with the final counterspeech. Moreover, participants were more likely to post the counterspeech co-created with CounterQuill on social media than ChatGPT.
6.2. Qualitative Findings of Learning with CounterQuill
6.2.1. Evolving Perceptions of Hate Speech: From Explicit to Implicit Forms
The learning session played a crucial role in shaping participants’ understanding of hate speech, particularly distinguishing between explicit and implicit forms. Many participants (19 — P1 - 13, P15 - P20) came into the study with assumptions about what does and does not constitute hate speech. While most (15 — P1 - P5, P7, P8, P10 - P13, P15 - P20) found it easy to recognize explicit hate speech, identifying implicit hate speech proved more challenging.
[…]
6.2.2. Shifting Perceptions: Recognizing the Lasting Impact of Implicit Hate Speech
Prior to the learning session, participants had varying levels of awareness about the impact of hate speech. Some, like P19, initially believed that implicit hate speech was less harmful than explicit hate speech, stating, I think (implicit hate speech) is less harmful (than explicit hate speech), it’s not direct, probably doesn’t hurt too many people..
However, the learning session, which included explaining the impact of different types of hate speech and testing participants’ knowledge, helped to create a unified understanding and changed many participants’ perceptions. After completing the session, most participants (17 — P1 - P5, P7, P8, P10 - P13, P15 - P20) recognized that despite its subtlety, implicit hate speech could have a more lasting and harmful impact than explicit hate speech. This shift in perspective is exemplified by P5’s reflections following the learning session:
@@ 6.2.2. Shifting Perceptions: Recognizing the Lasting Impact of Implicit Hate Speech
Participant 5: Before the learning session, I thought explicit hate speech was worse because it’s so obvious. But now I feel implicit [hate speech] has a much more lasting and harmful impact … The impact of implicit [hate speech] may be more rooted in society [be]cause only people [who belong to this] group can identify it as hate speech, [other people] may not think of it as hate speech.
Additionally, participants attributed the significant impact of implicit hate speech to several factors, including its pervasiveness (14 — P1, P4, P5, P7, P8, P10 - P13, P15-P19) and the difficulty of moderating it (8 — P2, P3, P10, P16 - P20).
6.2.3. Familiarizing with Counterspeech: From Awareness to Understanding
Our learning session introduced participants to the concept of counterspeech and its impact. Prior to the session, participants had varying levels of familiarity with counterspeech. Some participants (12 — P1 - P3, P8, P10 - P12, P14 - P17, P19) had a general understanding of counterspeech as a way to respond to hate speech. Two participants (P4, P13) had even posted responses to online hate speech prior to the session, but they were unaware that their behavior was considered counterspeech. Others (6 — P5, P6, P7, P9, P18, P20) were less familiar with the concept of counterspeech.
After completing the learning session, participants defined counterspeech as a respectful (18 — P2 - P12, P14 - P20) and non-escalating (8 — P1, P3, P6, P7, P9, P13, P18, P20) way to respond to hate speech. P4 provided a clear example of this understanding:
@@ 6.2.3. Familiarizing with Counterspeech: From Awareness to Understanding
Participant 4: I did respond to some hate [speech] before, but I never knew the term (of that behavior) … Counterspeech is a respectful way to address hate speech. It addresses (the targets are) human, and we should treat each other kindly … It’s a way to respond to hate [without] more hate, but [with an attempt to] get some understanding.
6.2.4. Empathy-Based Counterspeech: Recognizing Its Effectiveness
Our learning session also introduced participants to the concept of empathy-based counterspeech. Prior to the session, few participants (4 — P2, P4, P6, P18) were aware of this specific approach to counterspeech. However, after completing the learning session, many participants (19 — P1 - P12, P14 - P20) recognized the effectiveness of empathy-based counterspeech in countering hate content.
@@ 6.2.4. Empathy-Based Counterspeech: Recognizing Its Effectiveness
Overall, the learning session expanded participants’ understanding of counterspeech, particularly the value of empathy-based approaches in challenging hate speech effectively.
6.3. Qualitative Findings of Brainstorming with CounterQuill
6.3.1. Developing Hate Speech Identification Skills through Highlighting Practice
Our brainstorming session introduced participants to a color-coded highlighting practice, helping them identify the key components of hate speech: (1) Aspects of an individual or a group’s identity; (2) Action or perception that dehumanizes that person or group.
Prior to the highlighting practice, most participants (19 — P1 - 13, P15 - P20) had a general understanding of hate speech but struggled to articulate its specific components. As P11 mentioned, “I knew hate speech targeted certain groups and said bad things about them, but I couldn’t really break it down further than that.”.
However, after completing the highlighting practice, participants (17 — P1 - P13, P16, P17, P19, P20) found that the “the use of distinct colors (for targeted individuals or groups’ identities and dehumanizing actions) provided a clear visual breakdown of the hate speech components”, making it easier to grasp the core issues. P5 emphasized the benefits of this approach:
@@ 6.3.1. Developing Hate Speech Identification Skills through Highlighting Practice
Participant 5: The highlighting made it easy to get the main points [in the hate speech]. When I read the example, I got what was being said and the negative things [that targeted the group]. It showed what the hate speech was about and the specific bad things [being said].
Through practicing highlighting, most participants (16 — P1 - P7, P9 - P13, P16, P17, P19, P20) gained a clearer understanding of how to identify targeted identities and dehumanizing language in real-world contexts. Some participants (8 — P4, P7, P8, P12, P13, P15, P19, P20) mentioned that it was their first time recognizing these key components of hate speech. P12 shared how the practice enhanced their attention to specific words and stereotypical behaviors:
[…]
Participant 12: After completing the highlighting task, I would start looking for words that try to point at particular groups based on their ethnicity or sexual orientation. And, I would pay more attention to small actions that people take could be stereotyped. I’ll look out for when identifying the key things (in hate speech).
In addition to the hands-on practice of highlighting the targeted identity and dehumanizing action, most participants (15 — P1 - P7, P9 - P12, P15, P17, P19, P20) also described benefiting from the system’s highlighting feedback and comparison features (§4.2). The system-generated suggestions helped participants correct and refine their understanding of targeted identities and dehumanizing language. P13 described how the feedback helped her identify problematic aspects they had initially overlooked:
@@
Overall, the highlighting practice, coupled with the system’s feedback, effectively enhanced participants’ ability to identify and understand the key components of hate speech. As P6 summarized:
@@
Participant 6: … Doing this highlight practice has really increased my confidence when it comes to finding targeted groups and [dehumanizing] language [in hate speech]. Before I tried this, I probably missed some of the more subtle stuff … Everyone knows hate speech is bad, but after highlighting it will let you understand why it’s so toxic.
6.3.2. Developing Empathetic Reflection through Reflective Questions
Before engaging in brainstorming process, participants’ initial reactions to hate speech varied. Some (8 — P3, P5, P8, P11, P13, P14, P18, P19) mentioned that they would typically respond with anger or frustration, while others (6 — P1, P4, P6, P7, P10, P17) expressed uncertainty about how to counter hate speech effectively. However, many participants (16 —P1, P2, P4, P6, P7 - P11, P13, P14 - P18, P20) noted that answering the questions in the Brainstorming session changed their approach to responding to hate speech compared to their initial reactions. For example, P8 commented:

### v2

@@ 5.2. Study Protocol
Each 60-minute study session followed a within-subjects design. During sessions, participants wrote counterspeech using ChatGPT (20 minutes) and co-wrote with CounterQuill (40 minutes). Before the study, participants received a pre-study email with a tutorial video explaining CounterQuill’s features.
@@ 5.2.1. Conditions
We implemented two distinct conditions to evaluate the AI-assisted counterspeech writing, drawing on our findings in a preliminary survey (3.3) and previous research using GPT-based systems (Zhang et al., 2023; Reza et al., 2024; Lin et al., 2024) such as GPT-3.5 as baseline:
@@
  • CounterQuill: Participants used CounterQuill, which includes a learning session (help users understand hate speech and counterspeech), a brainstorming session (step-by-step counterspeech ideation support), and a writing session (co-writing with CounterQuill for counterspeech).
To minimize potential biases, we used a Latin Square experimental design (Ryan and Morgan, 2007) to ensure a balanced sequence of task topics and conditions for each study session: with half the participants starting with our system and the other half beginning with ChatGPT. To assess the potential impact of the system usage order on participants’ perceptions and the overall study outcomes, we conducted Welch’s Two-Sample t-tests that accounts for unequal variances between groups (Sakai, 2016), as detailed in Tables A1 and A2. The results confirmed that the sequence—whether participants began with ChatGPT or CounterQuill—did not significantly influence their performance or feedback on the system.
@@ 5.3. Study Procedure
At the beginning of each study session, the experimenters collected informed consent from the participants. For the user study, each participant was presented with 20 hate speech examples and asked to select one of them. The hate speech data was sourced from Ping’s (2024) research (Ping et al., 2024), which included a total of 400 hate speech instances evenly distributed across five hate speech themes: race, gender, sexual orientation, disability, and religion. To ensure a balanced representation of themes, each participant was randomly assigned 20 hate speech examples, with an average of 4 examples per theme. Appendix H lists the hate speech instances used by each of the 20 participants.
In the baseline condition, participants used ChatGPT to compose counterspeech responses to their selected hateful content, then completed a post-study questionnaire featuring NASA-TLX and 6 questions rating their counterspeech writing experience (Appendix D). In the CounterQuill condition, participants underwent a learning session about hate speech and counterspeech concepts, engaged in a structured brainstorming session for step-by-step ideation, and used the CounterQuill co-writing assistant before completing the same questionnaire. Both conditions concluded with a 20-minute semi-structured interview.
@@ 6. Result
6.1. Quantitative Evaluation
6.1.1. Comparative NASA-TLX Analysis
Table 3 summarizes the ratings for each dimension of the NASA Task Load Index (TLX) across two conditions: using ChatGPT (baseline) and using CounterQuill. We evaluated the differences between these conditions on each dimension using paired-sample $t$-tests (Mowery, 2011). To further contextualize these findings, Table 4 provides a comparison of the average time spent on counterspeech composition in each condition.
@@ 6.1.1. Comparative NASA-TLX Analysis
The analysis of the NASA-TLX dimensions revealed several notable differences between the two conditions. While Mental Demand was slightly higher with CounterQuill (though not statistically significant), Physical Demand, Temporal Demand, and Effort were higher when using CounterQuill. Performance was also significantly better with CounterQuill, showing a large effect size. Conversely, Frustration was the only dimension that was significantly lower when using CounterQuill compared to ChatGPT.
Task Average Time (min)
CounterQuill Learning 8.72 $\pm$ 1.02
CounterQuill Brainstorming 10.32 $\pm$ 1.55
CounterQuill Writing 10.48 $\pm$ 1.71
ChatGPT Writing 9.11 $\pm$ 1.06
In summary, CounterQuill demands greater effort, time, and energy than ChatGPT, but delivers higher perceived performance outcomes (as measured by NASA TLX ’Performance’), with participants reporting a 63.66% improvement. This stems not only from participants’ satisfaction with their co-written counterspeech but also from their strong performance in the post-learning session quiz, where 15 of 20 participants achieved perfect scores (91.25% average accuracy). Furthermore, CounterQuill reduced participants’ frustration levels by 51.96% compared to ChatGPT (as measured by ’Frustration’).
6.1.2. User Experience Ratings
Our study further investigated participants’ experiences and perceptions when using ChatGPT and CounterQuill for counterspeech writing. To gain deeper insights, we designed 6 custom questions (see Appendix D for detailed post-study questionnaire) that focused on specific aspects of the participants’ experience, such as confidence in identifying hate speech, effectiveness of the tools in providing guidance, and satisfaction with the final counter-speech. All questions were rated on a 7-point Likert scale.
@@ 6.1.2. User Experience Ratings
Confidence in Identifying Hate Speech 4.18$\pm$1.88 6.24$\pm$0.75 -4.280 ¡0.001*** [-3.079, -1.039]
Effectiveness in Brainstorming Ideas 3.24$\pm$1.44 6.35$\pm$0.70 -8.849 ¡0.001*** [-3.865, -2.371]
Confidence in Writing Counterspeech 3.29$\pm$1.83 6.18$\pm$0.81 -7.356 ¡0.001*** [-3.713, -2.052]
Engagement & Motivation While Using AI 3.59$\pm$1.73 6.00$\pm$0.87 -5.734 ¡0.001*** [-3.303, -1.520]
Satisfaction with the Final Counterspeech 4.00$\pm$1.58 6.47$\pm$0.51 -7.668 ¡0.001*** [-3.154, -1.788]
Willingness to Share Counterspeech Online 3.53$\pm$1.46 6.29$\pm$0.77 -10.444 ¡0.001*** [-3.326, -2.204]
Overall, CounterQuill demonstrated statistically superior outcomes compared to ChatGPT ($p$¡0.001). As shown in Table 5, key findings include:
  • Learning Session: Participants gained higher confidence in identifying online hate speech (HS Identification Confidence: 6.24 vs. 4.18).
  • Brainstorming Session: Participants rated CounterQuill as more effective for generating ideas (Brainstorming Effectiveness: 6.35 vs. 3.24) and reported stronger self-efficacy in ideating counterspeech (Self-Efficacy: 6.18 vs. 3.29).
  • Co-Writing Session: The collaborative process led to greater satisfaction with the final counterspeech (Satisfaction: 6.47 vs. 4.00) and increased willingness to share it online (Willingness to Post: 6.29 vs. 3.53).
6.2. Insights from the Learning Session
6.2.1. Learning About Online Hate Speech and Their Impact
The learning session enhanced participants’ ability to distinguish between explicit and implicit hate speech. While explicit hate speech was widely recognized, many participants (19 — P1–P13, P15–P20) initially struggled with identifying implicit hate speech. P7 noted, “Before the session, I knew explicit hate speech was bad, but I didn’t really think this (implicit hate speech) could also be hateful.”
After the session, participants (15 — P1–P5, P7, P8, P10–P13, P15–P20) expressed surprise at how subtle yet harmful implicit hate speech can be. P13 observed, “After finishing the learning session, I feel that because implicit hate speech is subtle, it requires deeper reflection to identify.” Similarly, P12 emphasized how implicit hate speech is often normalized in everyday discourse:
@@ 6.2.1. Learning About Online Hate Speech and Their Impact
6.2.2. Understanding Counterspeech and Their Impact
Before the session, participants had varying levels of familiarity with counterspeech. Some (14 — P1–P4, P8, P10–P13, P14–P17, P19) understood it as a general response to hate speech, while others (6 — P5, P6, P7, P9, P18, P20) were less familiar with the concept.
After completing the session, participants (18 — P2–P12, P14–P20) described counterspeech as a respectful and non-escalatory way to challenge hate speech. P4 provided a clear example:
@@ 6.2.2. Understanding Counterspeech and Their Impact
Participant 4: I did respond to some hate [speech] before, but I never knew the term (of that behavior) … Counterspeech is a respectful way to address hate speech. It addresses the targets [as] human, and we should treat each other kindly … It’s a way to respond to hate [without] more hate, but get some understanding.
6.2.3. Discovering Empathy-Based Responses
Few participants (4 — P2, P4, P8, P16) were previously aware of empathy-based counterspeech, but after the session, many (19 — P1–P12, P14–P20) recognized its effectiveness in de-escalating conflicts and fostering reflection.
@@ 6.2.3. Discovering Empathy-Based Responses
Overall, the learning session deepened participants’ understanding of hate speech and counterspeech, equipping them with more effective strategies for responding constructively.
6.3. Insights from the Brainstorming Session
6.3.1. Interactive Highlighting of Hate Speech Elements
Our brainstorming session introduced participants to a color-coded highlighting practice, helping them identify the key components of hate speech: (1) Aspects of an individual or a group’s identity; (2) Action or perception that dehumanizes the individual or group based on (1).
Prior to the highlighting practice, most participants (19 — P1 - 13, P15 - P20) had difficulty identifying specific components of hate speech. As P11 mentioned, “I knew hate speech targeted certain groups …, but I couldn’t really break it down further than that”. After completing the practice, participants (17 — P1–P13, P16, P17, P19, P20) reported that using different highlight colors helped them visually distinguish key components of hate speech, such as targeted identities and dehumanizing actions. P5 described the process:
@@ 6.3.1. Interactive Highlighting of Hate Speech Elements
Participant 5: The highlighting made it easy to get the main points [in the hate speech]. When I read the example, I got what was being said and the negative things [that targeted the group]. It showed what the hate speech was about and the specific bad things.
Many participants (15 — P1–P7, P9–P12, P15, P17, P19, P20) also found the system’s feedback on the users’ highlighted text using a side-by-side comparison of the correct highlights to be helpful. P13 explained how the feedback improved her awareness:
@@
Similiarly, P6 summarized how the highlighting practice boosted confidence in identifying hate speech:
@@
Participant 6: … Doing this highlight practice has really increased my confidence when it comes to finding targeted groups and dehumanizing language [in hate speech]. Before I tried this, I probably missed some of the more subtle stuff … Everyone knows hate speech is bad, but after highlighting it will let you understand why it’s so toxic.
6.3.2. Stimulating Reflective Ideation Through Guided Questions
Before engaging in brainstorming process, participants’ initial reactions to hate speech varied. Some (8 — P3, P5, P8, P11, P13, P14, P18, P19) mentioned that they would typically respond with anger or frustration, while others (6 — P1, P4, P6, P7, P10, P17) expressed uncertainty about how to counter hate speech effectively. However, many participants (16 —P1, P2, P4, P6, P7 - P11, P13, P14 - P18, P20) noted that answering the questions in the Brainstorming session led them to consider a more empathetic tone. For example, P8 commented: